OpenStudy (anonymous):

if y=sqrt(3x^2-3), determine dy/dt when x=4 and dx/dt=-1

OpenStudy (anonymous):

i know that the derivative of \[\sqrt{x}=\frac{ 1 }{2\sqrt{x} }\]

OpenStudy (anonymous):

@Jhannybean help? lol

jimthompson5910 (jim_thompson5910):

you're off to a good start, but now you have to use the chain rule

jimthompson5910 (jim_thompson5910):

what is the derivative of 3x^2-3 ?

OpenStudy (anonymous):

6x

jimthompson5910 (jim_thompson5910):

so using the chain rule gives you \[\Large y=\sqrt{3x^2-3}\] \[\Large \frac{dy}{dt}=\frac{1}{2\sqrt{3x^2-3}}*6x*\frac{dx}{dt}\]

jimthompson5910 (jim_thompson5910):

first you derive sqrt(3x^2-3) as if you were deriving sqrt(x) but then you have to derive the inside stuff 3x^2-3 and then you tack on dx/dt because when it comes to deriving the second inner stuff (the x) with respect to t, you get dx/dt

jimthompson5910 (jim_thompson5910):

the last thing to do is to plug in x=4 and dx/dt=-1 and simplify

OpenStudy (jhannybean):

You are not missing the derivative wrt x on the left hand side, @jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

you're thinking of \[\Large \frac{dy}{dx} = \frac{dy/dt}{dx/dt}\] \[\Large \frac{dy}{dx}*\frac{dx}{dt} = \frac{dy}{dt}\] at least I think you are

OpenStudy (jhannybean):

Oh, okay.

jimthompson5910 (jim_thompson5910):

but yeah that works too since the stuff in red is dy/dx \[\Large \frac{dy}{dt} = {\color{red}{\frac{dy}{dx}}}*\frac{dx}{dt}\] \[\Large \frac{dy}{dt} = {\color{red}{\frac{1}{2\sqrt{3x^2-3}}*6x}}*\frac{dx}{dt}\] so either way is fine

OpenStudy (anonymous):

the answer is \[\frac{ -4\sqrt{5} }{5}\] Thanks :)

jimthompson5910 (jim_thompson5910):

I'm getting the same thing. Nice job.